• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Rachek

    Log entry time 19:11:28 on January31,2002

    Entry number 81395

    keyword=Broken filter-board replaced

    Looking at gy:gx plot at calorimeter for high beam current runs
    one could have seen "dark spots" - areas with substantiously
    larger amount of events, although this is much less pronounced
    if high cluster energy cut is applied.
    Two most noticable regions are:

    * block #137
    * group of blocks which are summed in sum32 module #23

    Today I have checked these channels.

    * I found that there was a broken channel in filter board for
    block #137 . That resulted in an incorrect shape of pulse
    (it was abnormaly wide) at ADC for this channel and suppressed
    pulse height which was overcome by increasing high voltage at
    phototube. Another channel whose signal looks distorted is #135,
    which is also in this filter board.
    I have replaced the filter board by a spare one.
    High voltages were reduced for channels #137 & #135
    to get rouhgly the same laser pulse height as for other channels.
    Calibration for lead glass blocks whose signals passed through
    this filter-board should be checked/updated.

    * Nothing special was seen in sum32->discriminator #23 line,
    although a rate of double pulses at the output of this channel
    looks larger than for other channels, if low threshold is set.
    I changed discriminator channel (we have 8 spare channels in
    one of the modules). It seems that this new channel has smaller rate of
    double-pulses, although still a bit more than in other channels.

    FIGURE 1